### v2

# On-the-fly Surrogation for Complex Nonlinear Dynamics*

Abstract
High-fidelity models are essential for accurately capturing nonlinear system dynamics. However, simulation of these models is often computationally too expensive and, due to their complexity, they are not directly suitable for analysis, control design or real-time applications. Surrogate modelling techniques seek to construct simplified representations of these systems with minimal complexity, but adequate information on the dynamics given a simulation, analysis or synthesis objective at hand. Despite the widespread availability of system linearizations and the growing computational potential of autograd methods, there is no established approach that systematically exploits them to capture the underlying global nonlinear dynamics. This work proposes a novel surrogate modelling approach that can efficiently build a global representation of the dynamics on-the-fly from local system linearizations without ever explicitly computing a model. Using radial basis function interpolation and the second fundamental theorem of calculus, the surrogate model is only computed at its evaluation, enabling rapid computation for simulation and analysis and seamless incorporation of new linearization data. The efficiency and modelling capabilities of the method are demonstrated on simulation examples.
Index Terms:
Reduced order modelling, nonlinear systems, numerical algorithms.
I Introduction
Over the past century, linear time-invariant (LTI) dynamic models have been extensively used in engineering and industrial applications due to their simplicity and effectiveness [1]. Nonetheless, the increasing complexity and performance demands of modern systems often involve operational regimes where nonlinear dynamics inherent to physical processes play a more critical role, exceeding the capabilities of the LTI framework. Consequently, industrial applications require the construction of high-fidelity models, which are a detailed, precise representation of a physical system that closely replicates its actual behaviour. While these models provide high accuracy, they come at the cost of increased model complexity, in terms of model order, nonlinear dynamic relationships, dynamic coupling and computational load. This complexity often renders high-fidelity models not directly suitable for simulation, system analysis, controller design or real-time implementations.
These challenges manifested in the need for so-called surrogate models, which are low-complexity approximations/representations of all important information about the dynamics of the high-fidelity model for a given utilization objective. Various surrogate modelling frameworks have been developed, such as linear parameter-varying (LPV) systems [2, 3] and Koopman approaches [4, 5] which aim at giving a simplified, linear representation of the nonlinear system with respect to the input-output signal relations. However, these methods trade linearity for conservativeness and scheduling/state order of the resulting surrogate models, often making them difficult to use without further model reduction. Also applicability of LPV and Koopman modelling methods is limited to relatively moderate state order, and conversion to these forms often requires an explicit analytical form of the original model, or the methods are entirely data-based, requiring extensive trajectory-based representation of the system behaviour. Another widely used approach involves various types of neural networks (NN) or Gaussian process (GP) models to learn surrogate models from input-output simulation data [6, 7, 8]. However, these methods demand significant computational resources for the training process, their success largely depends on the generation of sufficiently informative data sets, and they are inherently approximative in their nature. Furthermore, the inclusion of new data requires retraining the obtained models, which often prevents real-time adaptation of these types of surrogates.
Additionally, computation of gradients of complicated nonlinear functions has become an inexpensive operation thanks to the recent development of automatic differentiation methods [9]. Consequently, obtaining system snapshots, i.e., local linear approximations of a nonlinear dynamical system at a specific point in operating space, is currently a cheap and reliable source of system information. Each snapshot contains a rich representation of the local system behaviour, and by collecting snapshots across the expected operating range, it is possible to generate an informative data set of the system behaviour. Despite these advantages, figuring out from these local linearizations an accurate global representation of system behaviour has been found challenging, often considered to be impossible, in the literature and has led to many local approximation methods since the dawn of gain scheduling [10, 11]. Hence, there is no established technique that systematically exploits local system snapshots to construct a surrogate model capable of capturing the entire global nonlinear dynamics. Existing approaches are either heavily approximative, rely on complex model conversion methods, or data-driven approximations that demand extensive computational resources and retraining upon the acquisition of new data. This gap motivates the need for a surrogate modelling framework that employs local linear information to reconstruct the global dynamics and allows the incorporation of new system observations without intensive computation processes.
In this work, we introduce a novel light-weight surrogate modelling approach that leverages system snapshots to construct a surrogate model whose computation is performed “on-the-fly”, i.e., the surrogate model is not assembled a priori, it is only computed at its evaluation, while it is learning the dynamics of the underlying nonlinear system in a global sense. To achieve this, we rely on two key contributions. First, we extend the radial basis functions (RBF)-based multidimensional scattered data interpolation approach [12, 13, 14, 15] to reconstruct the local linearized dynamics of the system from a set of snapshots from the system. Second, based on the second fundamental theorem of calculus (FTC) [16], we establish a one-to-one correspondence between the nonlinear dynamics and their local linearization. By combining these ingredients with numerical integration, a reliable approximation of the global dynamics at a time moment can be computationally efficiently obtained in the form of a surrogate model from the local dynamics. Compared to other surrogate modelling methods, our method offers two key advantages: (i) it provides a surrogate model that is not approximative in its nature, and we prove that it converges to the true nonlinear system representation as the number of system snapshots increases, and (ii) the construction of the surrogate model is computationally inexpensive and can be updated with new data seamlessly, as it is computed on-the-fly without requiring costly retraining or additional optimization steps.
The paper is structured as follows. In Section II, the surrogate modelling problem of nonlinear systems based on a set of local linearizations is formulated. For this problem, the proposed on-the-fly approach is presented in Section III, while in Section IV, the capabilities of the method are shown in simulation studies of a controlled Van der Pol oscillator and a mass-spring-damper system with nonlinear dynamics. Finally, in Section V, the main conclusions on the achieved results are drawn and possible further research directions are discussed.
Notation: The set of real numbers, integers, and positive integers are denoted as $\mathbb{R}$, $\mathbb{Z}$, and $\mathbb{Z}^{+}$, respectively. Let $a,b\in\mathbb{Z}^{+}$. The row-wise vectorization of a matrix $M\in\mathbb{R}^{a\times b}$ is given by $\operatorname{vec}(M)\in\mathbb{R}^{1\times ab}$, while its inverse (under row dimension $a$) is ${\operatorname{mat}_{a}}(\operatorname{vec}(M))=M$. Similarly, consider column-wise vectorization as $\operatorname{vec}^{\top}(M)$ and consecutive application of the operation in the form of $\operatorname{vec}^{\top}(M_{1},\ldots,M_{n})$. The gradient of the function $g:\mathbb{R}^{a}\to\mathbb{R}$ and the Jacobian of $g:\mathbb{R}^{a}\to\mathbb{R}^{b}$ w.r.t $z$ are denoted by $\nabla_{\mathrm{z}}(g)$ and $\mathrm{J}_{\mathrm{z}}(g)$, respectively.
II Problem definition
II-A System Description
Consider a time-invariant nonlinear system $\Sigma$ defined by the state-space representation
$\Sigma:\bigl{\{}\ \ \xi x(t)=f(x(t),u(t),\eta),$ (1)
where $t\in\mathbb{T}$ is time, $\xi$ is $\xi x(t)=\frac{d}{dt}x(t)$ in the continuous-time (CT) case with $\mathbb{T}=\mathbb{R}$ and $\xi x(t)=x(t+1)$ in the discrete-time case with $\mathbb{T}=\mathbb{Z}$. The signals $x(t)\in\mathbb{X}\subseteq\mathbb{R}^{n_{\mathrm{x}}}$ and $u(t)\in\mathbb{U}\subseteq\mathbb{R}^{n_{\mathrm{u}}}$ with $n_{\mathrm{x}},n_{\mathrm{u}}\in\mathbb{Z}^{+}$ are the states and inputs associated with (1), respectively. Moreover, $\eta\in\mathbb{H}\subseteq\mathbb{R}^{n_{\mathrm{\eta}}}$ with $n_{\mathrm{\eta}}\in\mathbb{Z}^{+}$ represents some physical parameters that affect the relation in (1). For the sake of simplicity, we consider the dynamics only in terms of the state evolution in (1); however, one can also consider output equations in the form of $y=h(x(t),u(t))$ with appropriate dimensions.
The function $f:\mathbb{R}^{n_{\mathrm{x}}}\times\mathbb{R}^{n_{\mathrm{u}}}\times\mathbb{R}^{n_{\mathrm{\eta}}}\to\mathbb{R}^{n_{\mathrm{x}}}$ is assumed to be at least one-time continuously differentiable, i.e., $f\in\mathcal{C}^{k}$ with $k\in\mathbb{Z}^{+}$, and, for any initial condition $x(t_{0})$ and any input trajectory $u(t)$, the solutions of (1) to be forward complete and unique for all $t\geq t_{0}$. Furthermore, let $\mathcal{X}\coloneq\{\underline{x}_{i}\leq x_{i}(t)\leq\overline{x}_{i}\}_{i=1}^{n_{\mathrm{x}}}\in\mathbb{X}$ denote the admissible state space, where $\underline{x}_{i}$ and $\overline{x}_{i}$ are the respective lower and upper bounds of the $i$-th state variable, which can be defined based on the operating range of the system. Likewise, let $\mathcal{U}:=\{\underline{u}_{i}\leq u_{i}(t)\leq\overline{u}_{i}\}_{i=1}^{n_{\mathrm{u}}}\in\mathbb{U}$ and :=$\mathcal{H}:=\{\underline{\eta}_{i}\leq\eta_{i}\leq\overline{\eta}_{i}\}_{i=1}^{n_{\mathrm{\eta}}}\in\mathbb{H}$ denote the admissible input and parameter spaces, respectively. Furthermore, let $z=\mathrm{vec}^{\top}(x,u,\eta)$ and also consider the admissible space of $\mathcal{Z}=\mathcal{X}\times\mathcal{U}\times\mathcal{H}$.
II-B Data collection
Let $\mathcal{L}_{f}:\mathbb{R}^{n_{\mathrm{x}}}\times\mathbb{R}^{n_{\mathrm{u}}}\times\mathbb{R}^{n_{\mathrm{\eta}}}\to\mathbb{R}^{n_{\mathrm{x}}\times(n_{\mathrm{x}}+n_{\mathrm{u}})}$ be the linearization operator of $f$, i.e., the operator that provides the linearized dynamics of (1) at a point $(x_{\ast},u_{\ast},\eta_{\ast})\in\mathbb{X}\times\mathbb{U}\times\mathbb{H}$:
$\mathcal{L}_{f}(x_{\ast},u_{\ast},\eta_{\ast})\coloneq\begin{bmatrix}\mathrm{J}_{\mathrm{x}}(f)&\mathrm{J}_{\mathrm{u}}(f)\end{bmatrix}(x_{\ast},u_{\ast},\eta_{\ast}).$ (2)
Given a set of local snapshots, i.e., linearizations of (1):
$\mathcal{M}$ :-$\coloneq\{\underbrace{\mathcal{L}_{f}(x^{(i)},u^{(i)},\eta^{(i)})}_{M_{i}}\}_{i=1}^{N},:$ (3)
where $\Theta\coloneq\{z^{(i)}\}_{i=1}^{N}$ with $z^{(i)}=\mathrm{vec}^{\top}(x^{(i)},u^{(i)},\eta^{(i)})$ is a set of $N\in\mathbb{Z}^{+}$ distinct observation points. The collection of available local information about (1) is denoted by
$\mathcal{D}\coloneq\{\Theta,\mathcal{M}\}.$ (4)
II-C Objective
Based on these considerations, our objective is to find a surrogate representation of the system in the form of
$\hat{\Sigma}:\bigl{\{}\ \ \xi\hat{x}(t)=\hat{f}(\hat{x}(t),u(t),\eta),$ (5)
where at any time moment $t$, i.e., for any value of $(\hat{x}(t),u(t),\eta)$, the true nonlinear function $f(\hat{x}(t),u(t),\eta)$ corresponding to the state increment $\xi\hat{x}(t)$ is approximated with the value $\hat{f}$ that is directly calculated based on the set of local linearizations $\mathcal{D}$ only.
Note that no analytic form of the mapping $\hat{f}:\mathbb{R}^{n_{\mathrm{x}}}\times\mathbb{R}^{n_{\mathrm{u}}}\times\mathbb{R}^{n_{\mathrm{\eta}}}\to\mathbb{R}^{n_{\mathrm{x}}}$ is required, only the true relation $f(\hat{x}(t),u(t),\eta)$ is aimed to be represented with (i) minimal approximation error and (ii) computationally efficient calculation of $\hat{f}(\hat{x}(t),u(t),\eta)$ based on $\mathcal{D}$.
III Methodology
III-A From local linearizations to global representation
As a first step to our objective, we establish a key relationship between the nonlinear state-space representation and its linearized dynamics. Building on the ideas in [16][17], and utilizing the FTC [18], we show that the the true nonlinear state equation given in (1) can be recovered completely from its linearizations given by $\mathcal{L}_{f}$ in (2).
Theorem 1
Given the nonlinear dynamics (1), the following identity holds true for any $(x_{\ast},u_{\ast})\in\mathbb{X}\times\mathbb{U}$:
$f(x,u,\eta)=f(x_{\ast},u_{\ast},\eta)+F(x,u,\eta)\begin{bmatrix}x-x_{\ast}\\ u-u_{\ast}\end{bmatrix},$ (6)
with
[…]
where $\lambda\in[0,1]$.
Proof.
For any nonlinear function $g_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$, $g_{i}\in\mathcal{C}^{1}$, scalar $\lambda\in[0,1]$ and point $z_{*}\in\mathbb{R}^{n}$, define the following auxiliary function $\bar{g}$ as follows
$\bar{g}_{i}(z;\lambda)=g_{i}(z_{\ast}+\lambda(z-z_{\ast})).$ (7)
Using the FTC and the identities $\bar{g}_{i}(1)=g_{i}(z)$, $\bar{g}_{i}(0)=g_{i}(z_{\ast})$, we obtain
[…]
Now, applying the chain rule of the integrand in (8), we obtain
$\frac{d\bar{g}_{i}}{d\lambda}(z;\lambda)=\Bigg{[}\Big{(}\nabla_{\mathrm{z}}g_{i}\Big{)}^{\top}\Big{(}z_{*}+\lambda(z-z_{*})\Big{)}\Bigg{]}(z-z_{\ast}).$ (9)
Substitution of (9) into (8) yields
[…]
Now, consider $g(z):=[g_{1}(z)\ \cdots\ g_{m}(z)]^{\top}$. Then, by stacking (10) column-wise for all $i\in[1,\cdots,m]$, $g$ is expressed as
[…]
Now, by applying (11) to the function $f$ in (1) leads to (6). ∎
Moreover, when $f(0,0,\eta)=0$, which corresponds to an equilibrium point in CT, Theorem 1 can be simplified.
Corollary 2
If $f(0,0,\eta)=0$ for a fixed $\eta\in\mathbb{P}$, then the right-hand side of (1) admits the following expression in terms of the linear dynamics:
$f(x,u,\eta)=F(x,u,\eta)\begin{bmatrix}x\\ u\end{bmatrix}$ (12)
where
[…]
Proof.
The proof is straightforward by substituting $(x_{\ast},u_{\ast},\eta)=(0,0,\eta)$ into (6) and using $f(0,0,\eta)=0$. ∎
The results from Corollary 2 allow the exact realization of the true nonlinear function $f$ from the linearization operator $\mathcal{L}_{f}$, provided $f(0,0,\eta)=0$. Note that when $f(0,0,\eta)\neq 0$, but $(0,0)\in\mathcal{X}\times\mathcal{U}$, we can perform a coordinate transformation to ensure this property.
However, application of Corollary 2 has a major drawback: it requires complete knowledge of the linearization of $f$, i.e., the operator $\mathcal{L}_{f}$, over the entire operating space. As in our problem setting, we only have a finite number of local snapshots of the system, we need to approximate $\mathcal{L}_{f}$ from the set of linearizations $\mathcal{D}$. This is what we tackle in the next subsection.
III-B Interpolation of local system dynamics
As mentioned previously, in our problem setting, we only have a finite number of local snapshots of the system. Specifically, we aim to leverage the fact that the system linearizations contained in $\mathcal{D}$ according to (4) provide an exact description of the original system (1) at the linearization points. Based on this reasoning, solving the surrogation problem amounts to finding a close approximation of $\mathcal{L}_{f}$ using the set $\mathcal{D}$ while introducing no error at the linearization points. This requirement is formulated as an interpolation problem as follows.
III-B1 Formulation of the interpolation problem
For a given set $\mathcal{D}$, find an interpolant $\mathcal{I}^{\mathcal{\mathcal{D}}}:\mathbb{R}^{n_{\mathrm{x}}}\times\mathbb{R}^{n_{\mathrm{u}}}\times\mathbb{R}^{n_{\mathrm{\eta}}}\to\mathbb{R}^{n_{\mathrm{x}}\times(n_{\mathrm{x}}+n_{\mathrm{u}})}$ such that the residual $r:\mathbb{R}^{n_{\mathrm{x}}}\times\mathbb{R}^{n_{\mathrm{u}}}\times\mathbb{R}^{n_{\mathrm{\eta}}}\to\mathbb{R}$ between $\mathcal{L}_{f}$ and $\mathcal{I}^{\mathcal{\mathcal{D}}}$, defined by
$r(z)\coloneq\|\mathcal{L}_{f}(z)-\mathcal{I}^{\mathcal{\mathcal{D}}}(z)\|_{2},$ (13)
satisfies the following conditions
$r(z)$ $=0\quad\text{for all}\quad z^{(i)}\in\Theta;$ (14a)
$r(z)$ $<\epsilon\quad\text{for all}\quad z^{(i)}\in\mathcal{Z}\setminus\Theta.$ (14b)
The problem of determining $\mathcal{I}^{\mathcal{\mathcal{D}}}$ subject to the requirements (14) is a multidimensional interpolation problem suggesting that regression-based methods, such as [19, 20, 21], are not suitable since they do not satisfy (14a).
Furthermore, one must not assume that the observation points contained in $\mathcal{D}$ are uniformly distributed. In reality, they are scattered instead.
III-B2 Multidimensional scattered interpolation with RBFs
Although there exist multiple alternatives to find $\mathcal{I}^{\mathcal{\mathcal{D}}}$ based on the given multidimensional scattered set $\mathcal{D}$, we propose the use of RBFs for this purpose, as this methodology has been proven to be an efficient and robust solution of the multidimensional scattered interpolation problem [22, 23, 24, 25]. The RBF-based interpolant that approximates $\mathcal{L}_{f}$ is given by
$\mathcal{I}_{\mathrm{RBF}}^{\mathcal{\mathcal{D}}}(z)=\operatorname{mat}_{n_{\mathrm{x}}}(\sum_{i=1}^{N}\alpha_{i}\phi(z-z^{(i)})),$ (15)
where $\phi$ is a RBF, i.e., a continuous function $\phi:[0,\infty)\to\mathbb{R}$ which is invariant under translation and rotation, $\alpha\in\mathbb{R}^{N\times n_{\mathrm{x}}(n_{\mathrm{x}}+n_{\mathrm{u}})}$, and the interpolation condition in (14a) is imposed as
$\mathcal{I}_{\mathrm{RBF}}^{\mathcal{\mathcal{D}}}(z^{(i)})=M_{i}\quad\text{for}\quad i\in\mathbb{I}_{1}^{N},$ (16)
where $\mathbb{I}_{a}^{b}\coloneq\{i\in\mathbb{Z}^{+}\mid a\leq i\leq b\}$ is an index set. Then, the coefficients $\alpha_{i}$ can be determined by formulating (16) as the following equation system:
$R\;\alpha=\gamma,$ (17a)
with
$[R]_{i,j}$ :-$\coloneq\phi(z^{(i)}-z^{(j)}),:$ $i,j\in\mathbb{I}_{1}^{N};$ (17b)
$[\gamma]_{i}$ :-$\coloneq\operatorname{vec}(M_{i}),:$ $i\in\mathbb{I}_{1}^{N},$ (17c)
where $R\in\mathbb{R}^{N\times N}$ and $\gamma\in\mathbb{R}^{N\times{n_{\mathrm{x}}(n_{\mathrm{x}}+n_{\mathrm{u}})}}$.
III-B3 Extension of the RBF interpolant with a polynomial tail
When variations of the local system dynamics are described by polynomial relationships of total degree less than $m$ the RBF-based interpolant (15) can be improved by adding a polynomial tail, as introduced in [14]. This polynomial tail represents the basis of $d$-variate polynomials of total degree at most $m-1$ which guarantees polynomial precision. Mathematically, the polynomial tail is denoted by $\mathcal{I}_{\mathrm{poly}}^{\mathcal{D}}$ and is defined as
$\mathcal{I}_{\mathrm{poly}}^{\mathcal{D}}(z)=\operatorname{mat}_{n_{\mathrm{x}}}(\sum_{j=1}^{Q}\beta_{j}q_{j}(z)),$ (18)
where $\beta\in\mathbb{R}^{Q\times n_{\mathrm{x}}(n_{\mathrm{x}}+n_{\mathrm{u}})}$ and $q_{1},\dots,q_{Q}$ is a basis for the $d$-variate polynomials of degree $\leq m-1$ with dimension
$Q=\frac{(m-1+d)!}{(m-1)!\,d!},\quad\text{with}\quad d=n_{\mathrm{x}}+n_{\mathrm{u}}+n_{\mathrm{\eta}}.$ (19)
In this case, the interpolant that approximates $\mathcal{L}_{f}$ becomes
$\mathcal{I}^{\mathcal{D}}(z)=\mathcal{I}_{\mathrm{RBF}}^{\mathcal{\mathcal{D}}}(z)+\mathcal{I}_{\mathrm{poly}}^{\mathcal{D}}(z),$ (20)
where now the interpolation condition (14a)
$\mathcal{I}^{\mathcal{D}}(z^{(i)})=M_{i},\quad i\in\mathbb{I}_{1}^{N},\$ (21)
is complemented with the side condition
$\sum_{i}^{N}\alpha_{i}q_{k}(z^{(i)})=0,\quad k\in\mathbb{I}_{1}^{Q},$ (22)
to cope with the additional degrees of freedom introduced by the polynomial tail.
III-B4 Determining coefficients of the interpolant
The coefficients $\alpha_{i}$ and $\beta_{j}$ related to $\mathcal{I}_{\mathrm{RBF}}^{\mathcal{\mathcal{D}}}$ and $\mathcal{I}_{\mathrm{poly}}^{\mathcal{D}}$ can now be determined by extending (17) with the side conditions (22), leading to:
$\begin{bmatrix}R&P\\ P^{\top}&0\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}=\begin{bmatrix}\gamma\\ 0\end{bmatrix},$ (23a)
with
$[P]_{i,j}$ :-$\coloneq q_{j}(z^{(i)}),\quad i\in\mathbb{I}_{1}^{N},\ j\in\mathbb{I}_{1}^{Q},:$ (23b)
where $P\in\mathbb{R}^{N\times Q}$, and $R$ and $\gamma$ are defined as in (17).
Now, we can show the uniqueness of solutions of (23) on the basis of [14, Chapter 8]. To this end, we first introduce the notion of conditionally positive definite (CPD) functions [14, Definition 8.1].
Definition 3
A continuous, even function $\phi:\mathbb{R}^{d}\to\mathbb{R}$ is said to be CPD of order $m$ if, for all pairwise distinct observation points $z_{1},\dots,z_{N}\in\mathbb{R}^{d}$ with $N\in\mathbb{Z}^{+}$ and all $\alpha\in\mathbb{R}^{N}\times\mathbb{R}^{n_{\mathrm{x}}\times(n_{\mathrm{x}}+n_{\mathrm{u}})}\setminus\{0\}$ that satisfy
$\sum_{i=1}^{N}\alpha_{i}q(z^{(i)})=0$ (24)
for all real-valued polynomials q of degree less than $m$, the following inequality holds that
$\sum_{i=1}^{N}\sum_{j=1}^{N}\alpha_{i}\alpha_{j}\phi(z^{(i)}-z^{(j)})>0.$ (25)
Then, using Definition 3, we have the following result that closely follows the treatment of [14, Theorem 8.21].
Theorem 4
Suppose that the function $\phi$ is a CPD of order $m$ and the columns of $P$ are linearly independent, where $P$ is given in (23b). Then, (23) has a unique solution.
Proof.
Suppose that $[\alpha^{\top}\ \beta^{\top}]^{\top}\in\mathbb{R}^{(N+Q)\times n_{\mathrm{x}}(n_{\mathrm{x}}+n_{\mathrm{u}})}$ is a solution of the homogeneous equation system in (23), i.e., with $\gamma=0$. Then, we have
$R\alpha+P\beta$ $=0;$ (26)
$P^{\top}\alpha$ $=0,$
where the second equation means that $\alpha$ satisfies the side condition (22). Now, pre-multiplying the first equation by $\alpha^{\top}$ gives $\alpha^{\top}R\alpha+\alpha^{\top}P\beta=\alpha^{\top}R\alpha+\left(P^{\top}\alpha\right)^{\top}\beta=\alpha^{\top}R\alpha=0$. As $\phi$ is CPD of order $m$, we can conclude that $\alpha=0$ and thus $P\beta=0$. Then, since the columns of $P$ are linearly independent, we can conclude that $\beta=0$. Therefore, (26) has trivial solutions only, meaning that the left-hand-side matrix of (23a) is full column rank, and thus (23) has a unique solution. ∎
Based on the results in Theorem 4, if the chosen RBF is CPD and the data points $z^{(i)}$ are assumed to be collected in a manner such that P in (23b) is full rank, (23) has a unique solution, which guarantees that, for a given $\mathcal{D}$, there exists a unique RBF interpolant extended with a unique polynomial tail in (20) that satisfies the interpolation condition (14a). One example of a CPD RBF function is the well-known Hardy multiquadric [14], which is CPD of (minimum) order $m=1$, given by
$\phi(z,c)=-\left(c^{2}+\|z\|_{2}^{2}\right)^{1/2},$ (27)
where $c>0$ is a constant shape parameter.
Now, for a given data set $\mathcal{D}$, we can use (15) to approximate the linearization operator $\mathcal{L}_{f}$ within the admissible space. Additionally, if the nonlinear system (1) is expected to contain polynomial relationships in its dynamics, we can use (20) instead to guarantee polynomial precision. In either case, the construction of the interpolant $\mathcal{I}^{\mathcal{D}}$ only involves an unconstrained linear system of equations, which can be efficiently solved with standard linear least squares methods [26]. In the next section, we detail how to utilize the proposed interpolant $\mathcal{I}^{\mathcal{D}}$ in combination with the results from Theorem 1 to achieve our objective of constructing our surrogate model on-the-fly .
III-C Finding the surrogate representation on-the-fly
Given the set $\mathcal{D}$, we can find a light-weight surrogate representation of (1) which is assembled on-the-fly. The first step of the process requires the pre-computation of the interpolant $\mathcal{I}^{\mathcal{D}}$ given by (20), which is executed in an offline fashion, as detailed in Algorithm 1. Second, assuming that $f(0,0,\eta)=0$, we use the results in Corollary 2 with $\mathcal{I}^{\mathcal{D}}$ to formulate the surrogate representation $\hat{\Sigma}$ as follows:
$\hat{\Sigma}\coloneq\bigl{\{}\xi\hat{x}=\hat{f}(\hat{x},u,\eta)=\hat{F}(\hat{x},u,\eta)\begin{bmatrix}\hat{x}\\ u\end{bmatrix},$ (28a)
where
[…]
and the integral in (28b) is left unevaluated. Lastly, this representation can be used to assemble the surrogate model on-the-fly, i.e., the surrogate model is only computed when it is required by solving the definite integral in (28b), as summarized in Algorithm 2. For this, it is necessary to invoke a numerical integration scheme to solve (28b) in a computationally efficient manner, as an analytic solution can be costly and often not computable.
Algorithm 1 Pre-computation of the interpolant (offline)
Input: Data set $\mathcal{D}$ of the form (4), RBF function $\phi(z,c)$, width parameter $c$, polynomial basis $q(z)$ of total degree $m$.
Output: Interpolant $\mathcal{I}^{\mathcal{D}}$ over $\mathcal{D}$.
1:Compute the matrices $R$, $P$ and $\gamma$, given by (17b), (23b), and (17c), respectively.
2:Find the coefficients $\alpha$ and $\beta$ by solving the system in (23).
3:Construct the interpolant $\mathcal{I}^{\mathcal{D}}$ given by (20).
4:Return $\mathcal{I}^{\mathcal{D}}$.
Algorithm 2 On-the-fly surrogation (online)
Input: Interpolant $\mathcal{I}^{\mathcal{\mathcal{D}}}$, value of $(\hat{x}(t),u(t),\eta)$ at a time moment $t$, numerical integration scheme.
Output: Approximated value of the state increment $\xi\hat{x}(t)$.
1:Formulate the surrogate representation $\hat{f}$ in (28a).
2:Compute $\xi\hat{x}(t)=\hat{f}(\hat{x}(t),u(t),\eta)$, solving the integral in (28b) with the numerical integration scheme.
3:Return $\xi\hat{x}(t)$.
The assembly of the surrogate model on-the-fly is a key feature of the proposed approach, as the inclusion of new data points only requires the computation of $\mathcal{I}^{\mathcal{D}}$ and thus is straightforward. This property is particularly advantageous for updating the surrogate model upon the acquisition of new system observations, which is a common scenario in real-world applications. In addition, the proposed approach guarantees that $\hat{f}$ converges to the true nonlinear system $f$ as the number of observations in $\mathcal{D}$ increases, leading to the following result.
Theorem 5
For a nonlinear system $\Sigma$ given by (1) where $f\in\mathcal{C}^{k}$ with $k\in\mathcal{Z}^{+}$, a set $\mathcal{D}=\{\Theta,\mathcal{M}\}$ given by (4) with $N\in\mathbb{Z}^{+}$ distinct linearizations, the interpolant $\mathcal{I}^{\mathcal{D}}$ given by (20) that approximates $\mathcal{L}_{f}$ in (2), and the surrogate $\hat{f}$ given by (28a), for a given point $(z_{\ast})\in\mathcal{Z}$, it holds that
$\|f(z_{\ast})-\hat{f}(z_{\ast})\|\to 0\quad$ (29)
as $N\to\infty$ with all linearization points considered to be distinct.
Proof.
The proof follows from the fact that $\mathcal{I}^{\mathcal{D}}$ satisfies the interpolation condition in (14a), therefore $\|\mathcal{L}_{f}-\mathcal{I}^{\mathcal{D}}\|\neq 0$ at any point $z\in\mathcal{Z}_{\Theta}^{\mathrm{c}}$, where $\mathcal{Z}_{\Theta}^{\mathrm{c}}$ denotes the relative complement of $\mathcal{Z}$ w.r.t. $\Theta$, i.e., $\mathcal{Z}_{\Theta}^{\mathrm{c}}\coloneq\mathcal{Z}\setminus\Theta$. Then, as $N\to\infty$, #$\#\mathcal{Z}_{\Theta}^{\mathrm{c}}\to 0$, where ##\## denotes the cardinality of the set. Thus, $\|\mathcal{L}_{f}-\mathcal{I}^{\mathcal{D}}\|\to 0$, and consequently, $\|f(z_{\ast})-\hat{f}(z_{\ast})\|\to 0$ by Theorem 1. ∎
III-D Optimization of the RBFs width parameter
In most cases, the RBFs are dependent on a width parameter $c$, for instance, see (27). The value of $c$ plays a major role in the achievable approximation performance of the interpolant $\mathcal{I}^{\mathcal{D}}$ in between observation points, and consequently, the accuracy of the surrogate $\hat{f}$. Therefore, based on the “leave-one-out” cross validation approach proposed in [15], we formulate the following optimization procedure to find $c$. For a given set of linearizations $\mathcal{D}=\{z^{(k)},M_{k}\}$, an RBF $\phi(z,c)$ and a polynomial basis $q$ of total degree $m$,
$c^{\ast}=\min_{c}\|[E_{1}\ldots E_{N}](c)\|_{p},$ (30)
with
$E_{k}(c)=M_{k}-I^{\mathcal{D}_{k}}(z^{(k)};c),\quad k\in\mathbb{I}_{1}^{N},$
where $I^{\mathcal{D}_{k}}$ denotes the interpolant given by (20) over the set $\mathcal{D}_{k}\coloneq\mathcal{D}\setminus\{z^{(k)}\}$, and $\|\cdot\|_{p}$ denotes the matrix induced $p$-norm. In particular, $p=\infty$ is in line with the objective (14b).
IV Examples
In this section, we present numerical examples to show the capabilities of the proposed on-the-fly surrogation approach. For this purpose, we provide two different examples. In the first example, we consider the Van der Pol oscillator, and our objective is to provide a clear and intuitive demonstration of the proposed method. In the second example, we consider an interconnection of mass-spring-damper (MSD) systems to show the scalability of the proposed approach. The full implementation of these examples is available at: https://gitlab.com/Javi-Olucha/cdc25-code-repo
IV-A Van der Pol oscillator
In this example, we consider the Van der Pol oscillator, whose dynamics are described as follows:
$\Sigma_{\mathrm{vdp}}\coloneq\bigl{\{}\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}x_{2}\\ -x_{1}-\eta x_{2}(1-x_{1}^{2})\end{bmatrix}+\begin{bmatrix}0\\ x_{1}u\end{bmatrix},$ (31)
where $\eta>0$ is a system parameter. Moreover, let $\Sigma_{\mathrm{vdp}}^{\tilde{\eta}}$ denote the Van der Pol oscillator with the parameter $\eta$ set at a fixed value $\tilde{\eta}>0$. Furthermore, we consider the infinite-horizon optimal control law given by $u=-x_{1}x_{2}$ for the performance objective $J(x,u)=x_{2}^{2}+u^{2}$ proposed by [27] that drives the system to the origin.
IV-A1 Reconstruction of the true nonlinear system from the linear dynamics
First, we demonstrate the best achievable performance from the methodology itself with the reconstruction of the global dynamics of the Van der Pol oscillator $\Sigma_{\mathrm{vdp}}^{0.5}$ based on perfect reconstruction of the underlying local linearizations. In this case, as the origin of $\Sigma_{\mathrm{vdp}}^{0.5}$ is an equilibrium, we can use the results in Corollary 2. Then, with the optimal control law $u$, we execute a CT simulation of $\Sigma_{\mathrm{vdp}}^{0.5}$ for each of the initial conditions $(-2,-2)$, $(-2,2)$, $(2,-2)$ and $(2,2)$.
Figure 1: State responses of the CT simulation of the Van der Pol oscillator with $\eta=0.5$, where ( ) is the true nonlinear system $\Sigma_{\mathrm{vdp}}^{0.5}$ and (   ) is the reconstruction of $\Sigma_{\mathrm{vdp}}^{0.5}$ using the results in Corollary 2 with a perfect reconstruction of the underlying local linearizations.
For the simulations, we use the Matlab in-built variable-step solver ode45 with the default parameters and step size, and simulate for 14 seconds. Next, we replicate the simulations using the same settings, but this time we use (12) instead of the true nonlinear system $\Sigma_{\mathrm{vdp}}^{0.5}$. For the integral in (12) we use the Matlab in-built function integral with the default tolerances, which computes a numerical approximation of the integral based on the global adaptive quadrature approach from [28]. For these simulations, ode45 requested an average of 21300 evaluations of the Jacobian of (31) with an average computation time of 0.16 seconds [^1] On a 64 bit laptop PC with 2.10 GHz 8-core CPU and 64 GB RAM. The simulation results, shown in Fig. 1, illustrate that the solutions of $\Sigma_{\mathrm{vdp}}^{0.5}$ and its reconstruction with Corollary 2 are identical, with an average root-mean-square error (RMSE) of $3.3\times 10^{-16}$, supporting the results in Subsection III-A.
IV-A2 On-the-fly surrogation of the Van der Pol oscillator
Second, we find a surrogate of $\Sigma_{\mathrm{vdp}}^{0.5}$ using the proposed on-the-fly surrogation method. For this, we define the following set of $N=9$ points:
$\small\Theta_{\mathrm{vdp}}^{1}=\left\{\!\begin{array}[]{cc}(-2,-2,-4),&(-2,2,4),\\ (2,-2,4),&(2,2,-4),\\ (-1.8,-1.8,-3.24),&(-1.8,1.8,3.24),\\ (1.8,-1.8,3.24),&(1.8,1.8,-3.24),\\ (0,0,0)\end{array}\!\right\}.$ (32)
Next, we construct the set $\mathcal{M}_{\mathrm{vdp}}^{1}$ by linearizing $\Sigma_{\mathrm{vdp}}^{0.5}$ at each point in $\Theta_{\mathrm{vdp}}^{1}$ and define the set $\mathcal{D}_{\mathrm{vdp}}^{1}=\{\Theta_{\mathrm{vdp}}^{1},\mathcal{M}_{\mathrm{vdp}}^{1}\}$. Then, we construct the interpolant $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{1}}$ with Algorithm 1, with the RBF $\phi$ given in (27) and the polynomial basis $\mathcal{I}_{\mathrm{poly}}(x,u)=\beta_{1}+\beta_{2}x_{1}+\beta_{3}x_{2}+\beta_{4}u$. The width parameter $c$ is determined by minimizing (30) with Bayesian optimization [29], since (30) is a non-smooth function. Note that other optimization methods suitable for non-smooth objectives can also be employed. Finally, we construct the surrogate representations $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{int}}$ and $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{cs38}}$ using $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{1}}$ and Algorithm 2 with two distinct numerical integration schemes. These two schemes are integral from Matlab, used for $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{int}}$, and the composite Simpson’s 3/8 rule (CS38) [30, Chapter 5.1] with 6 equidistant intervals, used for $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{cs38}}$.
Figure 2: State responses of the CT simulation of the Van der Pol oscillator with $\eta=0.5$, where ( ) is obtained with the true nonlinear system $\Sigma_{\mathrm{vdp}}^{0.5}$, ( ) with the surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{int}}$ and (   ) with the surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{cs38}}$.
Now, we compare the surrogates w.r.t. $\Sigma_{\mathrm{vdp}}^{0.5}$ by CT simulations, where the simulation and numerical integration settings for integral are set as described in the previous experiment IV-A1. For the simulations with $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{int}}$, ode45 requested an average of 20400 evaluations of $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{1}}$, leading to an average computation time and RMSE of 0.655 seconds and 0.1, respectively. For the simulations with $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{cs38}}$, ode45 requested an average of 816 evaluations of $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{1}}$, leading to an average computation time and RMSE of 0.028 seconds and 0.1, respectively. The results are shown in Fig. 2, where we observe that with only $N=9$ distinct linearization points, the obtained surrogates can approximate the behaviour of $\Sigma_{\mathrm{vdp}}^{0.5}$. Moreover, the use of CS38 drastically reduces the computation time, while the effects in numerical accuracy are negligible.
IV-A3 On-the-fly surrogation of the Van der Pol oscillator for a range of system parameter values
Third, we consider that the system parameter $\eta$ in (31) lies within the range $\eta\in[0.3,0.6]$. To capture the system dynamics for the different values of $\eta$, we define the set $\mathcal{D}_{\mathrm{vdp}}^{2}$, which contains $N=68$ system linearizations contained within the admissible spaces:
$\mathcal{X}_{\mathrm{vdp}}$ =$=\{\underline{x}_{i}=-2.2\leq x_{i}\leq\overline{x}_{i}=2.2\}_{i=1}^{2},$ (33)
$\mathcal{U}_{\mathrm{vdp}}$ =$=\{\underline{x}_{1}\underline{x}_{2}\leq u=-x_{1}x_{2}\leq-\overline{x}_{1}\overline{x}_{2}\},$
$\mathcal{H}_{\mathrm{vdp}}$ =$=\{\underline{\eta}=0.3\leq p\leq\overline{\eta}=0.6\}.$
Now, we construct the surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{2}}$ and compare in simulation with the true nonlinear system, using the same settings for the interpolant and the simulations described in the second experiment IV-A2, and CS38 with 6 equidistant intervals. In addition, for these simulations we set $\eta$ to different fixed values contained in (33) that are not included in $\Theta_{\mathrm{vdp}}^{2}$. In this case, ode45 also requested a maximum of 816 evaluations of the interpolant, leading to an average computation time and RMSE of 0.045 seconds and 0.784, respectively. These results are displayed in Fig. 3.
Figure 3: State responses of the CT simulation of the Van der Pol oscillator, where ( ) and ( ) are obtained with the true nonlinear systems ${\Sigma}_{\mathrm{vdp}}^{0.33}$ and ${\Sigma}_{\mathrm{vdp}}^{0.47}$, respectively, and (   ) and (   ) are obtained with the surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{2}}$ evaluated at $\eta=0.33$ and $\eta=0.47$, respectively.
IV-B Interconnection of MSD systems
In this example, we consider an interconnection of MSD systems depicted in Fig. 4, whose dynamics are given by:
Figure 4: Sketch of the interconnection of MSD systems.
$\Sigma_{\mathrm{msd}}\coloneq\left\{\begin{aligned} &m\ddot{x}_{1}=-F_{1}-F_{1,2};\\ &m\ddot{x}_{i}=-F_{i}-F_{i,\,i-1}-F_{i,\,i+1};\\ &m\ddot{x}_{\mathrm{Nm}}=-F_{\mathrm{Nm}}-F_{\mathrm{Nm},\,\mathrm{Nm}-1}+u,\end{aligned}\right.$ (34)
for $12\leq i\leq\mathrm{Nm}-$ 1, where $F_{i,\,j}$ and $F_{i}$ are the forces applied to the $i$-th mass by the adjacent masses and from the connection to the rigid wall, respectively, given by
$\begin{split}F_{i,j}&=k_{1}(x_{i}\!-\!x_{j})\!+\!b_{1}(x_{i}\!-\!x_{j})\!+\!(x_{i}\!-\!x_{j})^{3}\!+\!b_{2}(\dot{x}_{i}\!-\!\dot{x}_{j})^{3};\\ F_{i}&=k_{1}x_{i}+b_{1}\dot{x}_{i},\end{split}$
and the system parameters are chosen as the number of masses $\mathrm{Nm}=5$, $m=1$ kg, $k_{1}=0.5$ N/m, $b_{1}=1$ Ns/m, $b_{2}=2$ Ns/m, and the external force $u$ applied to the last mass is in Newtons. Further, we define the admissible spaces
$\mathcal{X}_{\mathrm{msd}}$ =$=\left\{\!\begin{array}[]{c}\{-2.2\leq x_{i}\leq=2.2\}_{i=1}^{5},\\ \{-1.5\leq x_{i}\leq=1.5\}_{i=6}^{10}\end{array}\!\right\};$ (35)
$\mathcal{U}_{\mathrm{msd}}$ =$=\{-1.5\leq u\leq=1.5\},$
and we construct the data set $\mathcal{D}_{\mathrm{msd}}$ with $N=100$ system linearizations at randomly drawn distinct points uniformly distributed within (35). Note that, as the respective state-order and input dimension of $\Sigma_{\mathrm{msd}}$ are $n_{\mathrm{x}}=10$ and $n_{\mathrm{u}}=1$, a regular grid containing only the vertices of the hyper-rectangle described by (35) requires $20482^{n_{\mathrm{x}}+n_{\mathrm{u}}}\!=\!$ 2048 system linearizations.
Similar as in Subsection IV-A2, we construct the interpolant $\mathcal{I}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$, where $\mathcal{D}_{\mathrm{msd}}$, the RBF $\phi$ in (27), the polynomial basis $\mathcal{I}_{\mathrm{poly}}(x,u)=\beta_{1}+\beta_{2}x_{1}+\dots\beta_{11}x_{10}+\beta_{12}u$, and the width parameter $c$ obtained by minimizing (30) with Bayesian optimization are the inputs to Algorithm 1. Then, we formulate the surrogate $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ according to Algorithm 2 using $\mathcal{I}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ and CS38 with 6 equidistant intervals.
Next, we test $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ by executing one thousand CT simulations, where the states corresponding to the mass velocities, i.e., $\{x_{i}\}_{i=6}^{10}$ are initialized at zero, and the states corresponding to the mass positions, i.e., $\{x_{i}\}_{i=1}^{5}$ are initialized at random values drawn within the admissible space defined in (35) with a uniform distribution. Moreover, the external force $u$ is set to $u(t)=0.7\sin(2\pi t)$, and each simulation is executed for 8 seconds using ode45 with the default parameters and step size. For these simulations, ode45 requested an average of 2352 evaluations of $\mathcal{I}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$, leading to an average computation time of 0.15 seconds. In Fig. 5
Figure 5: Simulated position (top) and velocity (bottom) trajectories of the $5$-th mass of the MSD interconnection, where ( ) and (   ) are obtained from the nonlinear system $\Sigma_{\mathrm{msd}}$ and the surrogate $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$, respectively.
Figure 6: Expected probability density distribution of the RMSE over one thousand simulations of $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ w.r.t. $\Sigma_{\mathrm{msd}}$, denoted as $e(\cdot,\cdot)$, where the expected probability density distribution is computed based on a normal kernel function and ( ) indicates the mean of the error.
we display the trajectory of the last mass obtained from some simulation results, where it is observed that the surrogate $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ approximates true nonlinear system $\Sigma_{\mathrm{msd}}$ with remarkable accuracy. Furthermore, we compute the RMSE for each state of the simulated $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ w.r.t. the simulated $\Sigma_{\mathrm{msd}}$. Then, we compute the expected probability density distribution of the RMSE data based on a normal kernel function [31], and besides some outliers with error values of up to $6\cdot 10^{-3}$, most of the error across all simulations is below $10^{-3}$, as shown in Fig. 6.
V Conclusions
We have presented a novel method that, based on a set of system linearizations, approximates the underlying nonlinear dynamics with a surrogate model that is not pre-computed, i.e., the surrogate model is computed “on-the-fly” when required. Compared to other methods, the main advantages of our approach are: (i) it approximates the global instead of the local underlying nonlinear dynamics, and we show that the proposed surrogate model converges to the true nonlinear system as the number of system linearizations increases; (ii) it avoids computationally intensive processes, such as the training of neural networks; (iii) the inclusion of new system observations only requires solving a linear least squares problem and is therefore straightforward; and (iv) the training data only consists of systems linearizations, which can be efficiently obtained from some simulation environments. In addition, we have demonstrated the global approximation capabilities of our method in various simulation experiments. For further research, we plan to investigate error estimates for the proposed surrogate modelling approach and the combination with state-order reduction techniques to enhance the scalability of our approach.